There are 5 places in a event. Each place can be won by a man or women. What are the outcomes?
Answer by dabanfield(803)

These are two outcomes for the first place (M or F). For each of these ways there are two outcomes for the second place and so on.
So for 5 events there are 2^5 = 2*2*2*2*2 = 32 ways.
You can list these:
MMMMM
FMMMM
MFMMM
FFMMM
etc.